# A volume-of-fluid model for biomass particle pyrolysis

### v1

@@ 4 Mass conservation
Simulations were performed for the three different cases: pure shrinking regime, constant volume (or fixed interface) and smooth function. Figure 2 shows the results considering $R_{0}=0.5$ m, $\epsilon_{0}=0.4$, $t_{end}=10$ s, $\rho_{s}=100$ kg m-3 and sum of formation rates $\sum_{i}^{NSS}\dot{\Omega}_{i}=-10$ kg m-3 s-1. 2(a) displays the profiles of the porosity along the radial direction. In the constant volume case ($Z=0$), it is possible to observe that the interface remains fixed while the porosity increases due to chemical reactions. On the contrary, in the pure shirking case, the porosity is identical to the initial snapshot, whereas the interface has regressed to preserve mass conservation. In the in-between situation, both phenomena occur simultaneously. The interface has regressed, but to a lesser extent, as part of the chemical reaction modified the porosity in the core region rather than solely contributing to particle shrinkage Furthermore, a gradient in the porosity field can be observed due to the different values of $Z$ in the two different regions. This result is also emphasised by 2(b), which shows a snapshot at the end time of the simulation for the smooth function case.
@@
Figure 3 shows the error in terms of the absolute difference between the numerical and the analytical solution. More in detail, 3(a) highlights two different behaviours in the order of convergence. In the case of pure porosity variation, second-order convergence is observed, whereas in the pure shrinking and smooth cases, a first-order convergence appears. The explanation for this is related to the fact that, with a fixed interface, the error is associated just with the initialisation of the volume fraction, i.e., the ability of the initialisation method to create a volume fraction that approximates a perfect circle. More accurate initialisation methods, such as the Vofi library [51], were able to reduce the error even further. In the other cases, the first-order error appears due to the accuracy of the method of the interface transport method of Weymouth and Yue [18] when applied to non-divergence-free flows. Nonetheless, convergence of both the solid and gas mass is achieved, highlighting the strength and correctness of the proposed solution method.
@@ 5 Model validation
Lastly, for all the following test cases, consider: i) the emissivity of wood for the interface ratiation to be $\alpha=0.9$ [52] ii) the solid particle to be initially at ambient temperature $T_{p}^{0}=298$ K and rapidly inserted into the hot environment so that the dynamics of this step can be neglected.
@@ 5.1 Isotropic sphere
300300400400500500600600700700800800020020040040060060080080010001000Refer to captioncorehalf-radiussurface
@@
Thermal conductivity $0.1937\times\omega_{\textit{wood}}+0.1405\times\omega_{\textit{char}}$ [W m-1 K-1]
@@
Figure 6 shows some zoomed snapshots of the temperature field at different times, together with arrows showcasing the gas flow field. Volatile release starts from the outer region and progresses through the particle, following the temperature evolution (Figure 6 and 6). Figure 6 also highlights the continuity of the gas velocity across the interface. Little numerical oscillations are observed in the interfacial cells due to the sharp variation of the Darcy drag term across it. Although subtle, the core temperature overshoot is also visible in Figure 6.
@@ 5.2 Wooden sphere in flow
Thermal conductivity $0.210\times\omega_{\textit{wood}}+0.071\times\omega_{\textit{char}}$ [W m-1 K-1]
@@
At the smaller diameter (10(a)), it can be observed that the model under-predicts the dependence of the final shrinking factor on the temperature. However, this behaviour seems to disappear for the larger particle, as seen in 10(b), showing a trend more in line with the model’s prediction. It is not clear why this tendency disappears for the bigger particles; further experimental measurements are necessary for a better understanding of this phenomenon. Nonetheless, the model correctly captures the experimental trend, predicting a sharp initial reduction in shape as the main components start to degrade. At lower temperatures, such as the case at $673$ K for the $2$ cm particle, the degradation is less pronounced in the first instants but increases once the particle reaches sufficiently high temperatures for all the main constituents to start degrading. As the particle reaches complete conversion, the shrinking stops, reaching a final diameter. The model overestimates the duration required for the shrinking factor to reach the final diameter and the corresponding value. However, the computed average error for the steady state value remains around $\approx 10\%$, which is reasonable given the simplified description of the shrinking problem and the high uncertainties and variability involved, including the correct biomass composition and possible effects of its anisotropy.
@@ 5.3 Anisotropic cylinder in forced convection
Figure 13 shows quantitatively the evolution of the particle shape and mass as pyrolysis progresses. 13(a) shows the time evolution of the shrinking factor, defined as the current over the initial dimension, along the radial $r$ and axial $z$ directions. As for the previous test cases, a delay can be noticed before the degradations start, due to the rise of temperature before reaching conditions for the pyrolysis reactions to be significant. Once the onset point is reached, the mass profile and particle dimensions sharply decrease. It can be seen that more shrinking is exhibited along the radial direction. This is mainly due to two factors: the presence of the hot wall at the top boundary, which increases heat transfer along the radial direction, and the anisotropic nature of the thermal conductivity. As this value is lower along the radial direction, heat transfer is hindered along this coordinate, leading to faster heating rates in the topmost cells. The model slightly underestimates the initial shrinking in both directions, probably due to some underestimation of the heat exchange in the first instants, related also to the sample insertion procedure. Nonetheless, the steady state values are accurately captured, considering the high variability and uncertainties in the properties of these systems. 13(b) shows that the model is in good agreement with the experimental measurements for the mass loss profile. Additionally, the final value of the remaining mass, a key indicator of the biochar yield of the process, is captured well.
To further validate our model, we compare the calculated interface heat exchange coefficients from our temperature field solution, in the first instants of the simulation, with those estimated by Gauthier et al. [57] from the experimental temperature profiles. For pure conduction (i.e. excluding radiation), our simulation yields $h_{cond}=12$ W m-2 K-1, compared to $15$ W m-2 K-1 from experiments. When radiation is included, the heat exchange coefficient increases to 46 W m-2 K-1 in our numerical simulation, compared to 42 W m-2 K-1 in the experimental study. The discrepancy appears because Gauthier et al. [57] uses a linearising function to estimate the radiative contribution. This comparison demonstrates good agreement between our model and experiments, while also confirming the significant role of radiation in the particle’s heating behaviour and the model’s capacity to serve as an exploration tool for sub-grid correlations.
@@ 6 Conclusion
This work proposes a novel methodology for describing the evolution of biomass particle pyrolysis. The framework relaxes the constraints of a fixed particle interface and the need for heat and mass transfer coefficient correlations, which are often employed as tunable parameters to match experimental measurements. Instead, these effects are computed directly by solving the surrounding flow field, while accounting for an arbitrarily moving interface. To achieve this, a Volume-Of-Fluid model was modified to describe the evolution of a porous phase, incorporating pyrolysis chemical kinetics, Darcy–Forchheimer contributions, and anisotropic transport properties. The model demonstrated first-order convergence in mass conservation. Comparisons with experimental data showed that the model captures temperature and mass loss profiles accurately, while the release of volatiles follows the correct experimental trends. The proposed approach for distributing porosity changes and volume variation yields reasonable trends, but also highlights a lack of fundamental understanding regarding the evolution of the internal particle structure. The model can be used to predict yields and characteristic degradation times, aiding the development of pyrolysis processes even under complex fluid dynamic conditions. The model was developed within the open-source code Basilisk, and released in open-source form, in order to enhance the reproducibility of the results, and allowing improvements from the community.

### v2

@@ 4.2 Mass conservation
Simulations were performed for the three different cases: pure shrinking regime, constant volume (or fixed interface) and smooth function. Figure 4 shows the results considering $R_{0}=0.5$ m, $\epsilon_{0}=0.4$, $t_{end}=10$ s, $\rho_{s}=100$ kg m-3 and sum of formation rates $\sum_{i}^{NSS}\dot{\Omega}_{i}=-10$ kg m-3 s-1. 4(a) displays the profiles of the porosity along the radial direction $r$. In the constant volume case ($Z=0$), it is possible to observe that the interface remains fixed while the porosity increases due to chemical reactions. On the contrary, in the pure shirking case, the porosity is identical to the initial snapshot, whereas the interface has regressed to preserve mass conservation. In the in-between situation, both phenomena occur simultaneously. The interface has regressed, but to a lesser extent, as part of the chemical reaction modified the porosity in the core region rather than solely contributing to particle shrinkage. Furthermore, a gradient in the porosity field can be observed due to the different values of $Z$ in the two different regions. This result is also emphasised by 4(b), which shows a snapshot at the end time of the simulation for the smooth function case.
@@
Figure 5 shows the error in terms of the absolute difference between the numerical and the analytical solution. More in detail, 5(a) highlights two different behaviours in the order of convergence. In the case of pure porosity variation, second-order convergence is observed, whereas in the pure shrinking and smooth cases, a first-order convergence appears. The explanation for this is related to the fact that, with a fixed interface, the error is associated just with the initialisation of the volume fraction, i.e., the ability of the initialisation method to create a volume fraction that approximates a perfect circle. More accurate initialisation methods, such as the Vofi library [11], were able to reduce the error even further. In the other cases, the first-order error appears due to the accuracy of the method of the interface transport method of Weymouth and Yue [69] when applied to non-divergence-free flows. Nonetheless, convergence of both the solid and gas mass is achieved, highlighting the strength and correctness of the proposed solution method.
@@ 5 Model validation
The proposed mathematical and numerical formulation is independent of the number and shape of biomass particles present. The model can be used to study arbitrary combinations of the latter, as all information is captured in the volume fraction scalar field. However, experimental measurements of interactions among multiple particles under simplified conditions are, to the author’s knowledge, unavailable. Therefore, the proposed model is validated on individual particles, for which data are widely available, and will be used in the future to study the effects of such interactions on yields and heating profiles. The possibility of including multiple particles underscores the model’s capacity to serve as a sub-grid model finder for larger-scale computational tools.
@@
A grid independence study was performed for each of the following cases, so the results presented in the next part are to be considered as not affected by grid resolution. The following results at different grid refinement levels are made available as supplementary materials.
The presented experiment and simulations are conducted under conditions in which the isothermal assumption is invalid, as demonstrated by the internal temperature profiles. This is particularly accentuated due to additional thermal effects arising from pyrolysis reactions and water evaporation, further justifying the need for a spatially resolved multidimensional model.
Lastly, for all the following test cases, consider: i) the emissivity of wood for the interface radiation to be $\alpha=0.9$ [43] ii) the solid particle to be initially at ambient temperature $T_{p}^{0}=298$ K and rapidly inserted into the hot environment so that the dynamics of this step can be neglected.
@@ 5.1 Isotropic sphere
300300400400500500600600700700800800020020040040060060080080010001000Refer to captionCoreHalf radiusSurface
@@
Thermal conductivity $0.1937\times\omega_{\textit{wood}}+0.1405\times\omega_{\textit{char}}+\lambda_{g}\epsilon$ [16] [W m-1 K-1]
@@
Figure 8 shows some zoomed snapshots of the temperature field at different times, together with arrows showcasing the gas flow field. Volatile release starts from the outer region and progresses through the particle, following the temperature evolution (Figure 8 and 8). Figure 8 also highlights the continuity of the gas velocity across the interface. Little numerical oscillations are observed in the interfacial cells due to the sharp variation of the Darcy drag term across it. Although subtle, the core temperature overshoot is also visible in Figure 8.
@@ 5.2 Wooden sphere in flow
Thermal conductivity $0.210\times\omega_{\textit{wood}}+0.071\times\omega_{\textit{char}}+\lambda_{g}\epsilon_{g}$ [23] [W m-1 K-1]
@@
At the smaller diameter (12(a)), it can be observed that the model under-predicts the dependence of the final shrinking factor on the temperature. However, this behaviour seems to disappear for the larger particle, as seen in 12(b), showing a trend more in line with the model’s prediction. It is not clear why this tendency disappears for the bigger particles; further experimental measurements are necessary for a better understanding of this phenomenon. Nonetheless, the model correctly captures the experimental trend, predicting a sharp initial reduction in shape as the main components start to degrade. At lower temperatures, such as the case at $673$ K for the $2$ cm particle, the degradation is less pronounced in the first instants but increases once the particle reaches sufficiently high temperatures for all the main constituents to start degrading. As the particle reaches complete conversion, the shrinking stops, reaching a final diameter. The model overestimates the duration required for the shrinking factor to reach the final diameter and the corresponding value. However, the computed average error for the steady state value remains around $\approx 10\%$, which is reasonable given the simplified description of the shrinking problem and the high uncertainties and variability involved, including the correct biomass composition and possible effects of its anisotropy.
@@ 5.3 Anisotropic cylinder in forced convection
Figure 15 shows quantitatively the evolution of the particle shape and mass as pyrolysis progresses. 15(a) shows the time evolution of the shrinking factor, defined as the current over the initial dimension, along the radial $r$ and axial $z$ directions. As for the previous test cases, a delay can be noticed before the degradations start, due to the rise of temperature before reaching conditions for the pyrolysis reactions to be significant. Once the onset point is reached, the mass profile and particle dimensions sharply decrease. It can be seen that more shrinking is exhibited along the radial direction. This is mainly due to two factors: the presence of the hot wall at the top boundary, which increases heat transfer along the radial direction, and the anisotropic nature of the thermal conductivity. As this value is lower along the radial direction, heat transfer is hindered along this coordinate, leading to faster heating rates in the topmost cells. The model slightly underestimates the initial shrinking in both directions, probably due to some underestimation of the heat exchange in the first instants, related also to the sample insertion procedure. Nonetheless, the steady state values are accurately captured, considering the high variability and uncertainties in the properties of these systems. 15(b) shows that the model is in good agreement with the experimental measurements for the mass loss profile. Additionally, the final value of the remaining mass, a key indicator of the biochar yield of the process, is captured well.
To further validate our model, we compare the calculated interface heat exchange coefficients from our temperature field solution, in the first instants of the simulation, with those estimated by Gauthier et al. [28] from the experimental temperature profiles. For pure convection (i.e. excluding radiation), our simulation yields $h_{cond}=12$ W m-2 K-1, compared to $15$ W m-2 K-1 from experiments. When radiation is included, the heat exchange coefficient increases to 46 W m-2 K-1 in our numerical simulation, compared to 42 W m-2 K-1 in the experimental study. The discrepancy appears because Gauthier et al. [28] uses a linearising function to estimate the radiative contribution. This comparison demonstrates good agreement between our model and experiments, while also confirming the significant role of radiation in the particle’s heating behaviour and the model’s capacity to serve as an exploration tool for sub-grid correlations.
@@ 6 Conclusion
This work proposes a novel methodology for describing the evolution of biomass particle pyrolysis. The framework relaxes the constraints of a fixed particle interface and the need for heat and mass transfer coefficient correlations, which are often employed as tunable parameters to match experimental measurements. Instead, these effects are computed directly by solving the surrounding flow field, while accounting for an arbitrarily moving interface. To achieve this, a Volume-Of-Fluid model was modified to describe the evolution of a porous phase, incorporating pyrolysis chemical kinetics, Darcy–Forchheimer contributions, and anisotropic transport properties. The model showed first-order convergence in particle mass conservation , due to the non-divergence-free velocity and the VOF interface transport. Rise to second order is observed for stationary interfaces, where volume fraction initialisation controls the convergence rate. Comparisons with experimental data showed that the model captures temperature and mass loss profiles with decent accuracy, while the release of volatiles follows the experimental trends. The proposed approach for distributing porosity changes and volume variation yields reasonable trends, but also highlights a lack of fundamental understanding regarding the evolution of the internal particle structure. The model can be used to predict yields and characteristic degradation times, aiding the development of pyrolysis processes even under complex fluid dynamic conditions. While providing insights into the degradation behaviour of biomass, the model can also serve as a tool for multi-scale modelling of biomass conversion processes. In particular, this model enables detailed studies of single particles of arbitrary geometry under controlled conditions. Additionally, the proposed formulation enables the study of interactions among multiple particles. From these studies, drag, heat, and mass transfer coefficients characterising the systems can be obtained, as well as information on the interplay between porosity and volume changes. This information can be transformed into sub-grid-scale correlations, which are necessary for laboratory- or industrial-scale reactor simulations. Euler-Lagrange [39] or Euler-Euler [62] models, which do not resolve the interface between the particle and the gas phase, usually employ these correlations. This hierarchical strategy for simulating the biomass gasification process will ultimately enable optimisation of gasification and pyrolysis processes with respect to conversion, yields, operating conditions, efficiency, and emissions. The model was developed within the open-source code Basilisk, and released in open-source form, in order to enhance the reproducibility of the results, and allowing improvements from the community.